OpenStudy (anonymous):

What is m∠CDG? Right angle with intersecting line forming a 35 degree angle - assessment graphic A. 35° B. 55° C. 90° D. 125°

OpenStudy (anonymous):

I think the answer is B.

OpenStudy (anonymous):

@tkhunny

OpenStudy (phi):

a right angle is 90 degrees. the two angles <CDG + <GDF add up to 90

OpenStudy (phi):

so 35 + <CDG = 90

OpenStudy (anonymous):

35+55=90

OpenStudy (phi):

you can test each answer, or you can use "algebra" <CDG= 90 - 35

OpenStudy (anonymous):

90-35=55

OpenStudy (anonymous):

oh by the way I am not new I just created myself a new profile , my old profile name was rainbow_rocks03

OpenStudy (phi):

in case you do not know, if you are able to write <CDG + 35 is 90 or <CDG+ 35= 90 then "add -35" to both sides like this <CDG + 35 + -35 = 90 + -35 and simplify on the left side +35-35 is 0 and <CDG + 0 is <CDG on the right side you get 55 <CDG = 55

OpenStudy (anonymous):

So my answer is 55 ?

OpenStudy (phi):

yes

OpenStudy (anonymous):

YAY! I got it right!

OpenStudy (anonymous):

I will go back to my old profile actually so I will fan you under the name of rainbow_rocks. :)